OpenStudy (freemap):

Luis can spend up to $240 buying men’s and women’s shirts for his club. Men’s shirts cost $8 each and women’s shirts cost $12 each. He needs to buy at least 3 of each type of shirt. Also, the women’s shirts must be at least twice the number of men’s shirts. Which combinations of men’s and women’s shirts can Luis buy? Choose all answers that are correct. A. 4 men’s shirts and 9 women’s shirts B. 5 men’s shirts and 8 women’s shirts C. 6 men’s shirts and 15 women’s shirts D. 7 men’s shirts and 17 women’s shirts

OpenStudy (freemap):

@AdityaOS

OpenStudy (freemap):

I have D for one answer

OpenStudy (freemap):

I think

OpenStudy (anonymous):

B option incorrect since women's shirt is not twice of men's shirt

OpenStudy (anonymous):

for option A 12*9 = 108 and 8*4 = 32 108+32=140 which is less than 240 so it is also correct

OpenStudy (anonymous):

as for c option for women => 12*15=180 and for men 8*6 = 48 180+48 = 228 again less than 240 so correct

OpenStudy (anonymous):

no for option D => women => 12*17= 204 for man => 8*7 = 56 here 204 + 56 = 260 which is greater than 240 hence wrong option

OpenStudy (freemap):

Thank you. I really appreciate you helping me. Thanks again

OpenStudy (anonymous):

it's ok no problem